OpenStudy (anonymous):

-1+19w=11w+23 9k-26=6k-8 2y+7=y Show all work please

OpenStudy (anonymous):

what i dont understand tht could u try seperating them and show work please

OpenStudy (anonymous):

they are 3 different problems just to let yall knw

OpenStudy (diyadiya):

\[-1+19w=11w+23\] \[8w=24 \] \[w=\frac{24}{8}\] \[w=3\]

OpenStudy (diyadiya):

if you can try the next one , I'll tell you if its right or wrong

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

ok i dont understand how u got 8w=24

OpenStudy (diyadiya):

Oh i skipped a step there

OpenStudy (diyadiya):

\[19w-11w=23+1\]

OpenStudy (anonymous):

ok i dont understand how ur doing this can u draw it and show me

OpenStudy (diyadiya):

ok

OpenStudy (diyadiya):

your question is -1+19w=11w+23 first subtract 11w from bothsides -1+19w -11w =11w+23-11w -1+19w-11w=23 did you understand this ?

OpenStudy (diyadiya):

we are doing this to bring the terms with variable on one side

OpenStudy (anonymous):

ok so now i have -1+8w=23 now what

OpenStudy (diyadiya):

now add 1 on both sides

OpenStudy (diyadiya):

so that we get the term with variable on one side and the constants on other side

OpenStudy (diyadiya):

-1+8w+1=23+1

OpenStudy (diyadiya):

8w=24

OpenStudy (anonymous):

ok and then w=3 now lets work on the next one

OpenStudy (diyadiya):

Now divide bothe sides with 8 \[\frac{8w}{8}=\frac{24}{8}\] \[w=\frac{24}{8}\] \[w=3\]

OpenStudy (diyadiya):

ok Try it

OpenStudy (anonymous):

ok so 9k - 26 = 6k - 8 would u subtract both sides by 6k

OpenStudy (diyadiya):

Yes ! :)

OpenStudy (anonymous):

so it would be 3k-26=8

OpenStudy (diyadiya):

no 3k-26=-8

OpenStudy (anonymous):

ok now what im not sure

OpenStudy (diyadiya):

you can add 26 on both sides

OpenStudy (anonymous):

ok then 3k=34 ?

OpenStudy (anonymous):

??? or is tht wrong

OpenStudy (diyadiya):

3k=26-8

OpenStudy (anonymous):

so k=6

OpenStudy (diyadiya):

yes

OpenStudy (anonymous):

ok next one hey btw i have more problems if you will stay on this thing and work them out with me thanks

OpenStudy (anonymous):

2y+7=y

OpenStudy (diyadiya):

hm you can post it sorry im li'l busy right now

OpenStudy (diyadiya):

y=-7

OpenStudy (anonymous):

i need it to be worked out not just the answer

OpenStudy (anonymous):

how would u work this out